An economy has the following total transactions input-output matrix:

Agriculture Manufacturing Energy Services
Agriculture 1.3 .4 .3 .4
Manufacturing .4 1.5 .4 .4
Energy .3 .5 1.2 .6
Services .5 .5 .6 1.2

(a) If final demand (say exports) of energy products rises by $ 300 billion, what will be the increase in output in each industry? In GDP?

(b) Calculate the GDEP multipliers for agriculture and manufacturing.

(c) In using the input-output model, what did you assume about input and output prices?
Answer #1

a) If energy products increase by 300 billion dollars then,

Output of agriculture will increase by 90 billion dollars (300*0.3).

Output of manufacturing will increase by 120 billion dollars (300*0.4).

Output of energy will increase by 360 billion dollars (300*1.2).

Output of services will increase by 180 billion dollars (300*0.6).

b) gdep multiplier is equal to the vertical summation of each column.

Agriculture multiplier =1.3+0.4+0.3+0.5=2.5

Manufacturing multiplier=0.4+1.5+0.5+0.5=2.9

c) Input and output prices are specified in advance and assumed to be not changing and it is also assumed that there will be perfect competition in the economy.